# AB-Training: A Communication-Efficient Approach for Distributed Low-Rank Learning

### v1

@@ 4.1. Computational Environment
4.2. Datasets and Models
We used the ImageNet-2012 dataset, which contains 1.2 million images, for our scaling experiments. Basic image augmentation was applied, including normalization, random resizing, random cropping, and a random flip. We trained two models on this dataset: ResNet-50 (He et al., 2016) and Vision Transformer B/16 (ViT) (Dosovitskiy et al., 2021), chosen for their widespread use and distinct architectures.
We also trained VGG16 on the CIFAR-10 dataset, which contains 50,000 images. Basic image augmentation was applied in the same way as for the ImageNet experiments.
All models were trained with the AdamW (Loshchilov and Hutter, 2019) optimizer. Other training hyperparameters can be found in Appendix 1.
4.3. Hyperparameter Considerations
Careful hyperparameter tuning is essential for balancing the communication reduction achieved through our method with potential impacts on training stability and accuracy. Key hyperparameters (HPs) include the warmup and full-rank rebound phase durations, the number of AB training iterations, the AB decomposition rank-reduction parameter, and the frequency of SVD and synchronization steps. Additionally, the latter will influence the trade-offs between regularization and communication efficiency.
Our hyperparameter search on CIFAR-100 (Krizhevsky, 2009) with a minified Vision Transformer model (patch size of eight instead of 16, six heads instead of twelve, depth of six instead of twelve) using Propulate (Taubert et al., 2023) suggested the following guidelines:
@@ 4.3. Hyperparameter Considerations
4.4. Constant Local Batch Size Scaling
@@
Figure 3 shows the highest top 1 accuracy achieved by each model during training. Figure 4 shows the lowest binary cross-entropy loss values achieved across all epochs. The compression values for the trained models are shown in Figure 5.
@@
Figure 3. Highest top 1 accuracy measurements for each training run on ImageNet-2012 for two network architectures with a constant local batch size of 256. Global batch sizes range from 2,048 to 32,768 in powers of 2. Error are plotted, though not always visible.
[…]
Highest top 1 accuracy measurements for training runs on ImageNet-2012 for two network architectures with a constant local batch size of 256. Global batch sizes range from 2,048 to 32,768 in powers of 2.
@@
Figure 4. Lowest binary cross-entropy (the loss function used during training) measurements during each training run on ImageNet-2012 for two network architectures with a constant local batch size of 256. Global batch sizes range from 2,048 to 32,768 in powers of 2. Error are plotted, though not always visible.
[…]
Lowest binary cross-entropy (the loss function used during training) measurements for training runs on ImageNet-2012 for two network architectures with a constant local batch size of 256. Global batch sizes range from 2,048 to 32,768 in powers of 2. AB training is shown to have a regularization effect for both ResNet-50 and the Vision Transformer. Large batch effects are also visible as the loss values get worse (larger) as the global batch size increases.
Figure 5. Average compression ratios for AB trained models against the traditional data parallel baseline trained models on ImageNet-2012 for two network architectures with a constant local batch size of 256.
[…]
Compression ratios for AB trained models against the traditional data parallel baseline trained models on ImageNet-2012 for two network architectures with a constant local batch size of 256.
4.5. Constant Global Batch Size Scaling
In this experiment, we maintain a fixed global batch size while scaling compute resources. This contrasts with the standard approach of keeping local batch sizes constant, emphasizing the communication overhead reductions achieved by AB training. With a fixed global batch size of 4,096, a four-node run would have a local batch size of 256, whereas an eight-node run would have a local batch size of 128.
This experiment directly quantifies the communication benefits of our method in scenarios where local batch sizes become very small due to scaling. It also provides insight into the ability of independent training groups to learn meaningful representations when the batch size used for independent training is much smaller than the global batch size. Results for these runs are shown in Figures 6 and 7.
(a) Vision Transformer B/16
(b) ResNet-50
@@ 4.5. Constant Global Batch Size Scaling
Figure 6. Highest top 1 accuracy measurements for each training run on ImageNet-2012 for two network architectures with a constant global batch size of 4,096. Error are plotted, though not always visible.
[…]
Highest top 1 accuracy measurements for training runs on ImageNet-2012 for two network architectures with a constant global batch size of 4,096.
@@
Figure 7. Compression ratios for AB trained models against the traditional data parallel baseline trained models on ImageNet-2012 for two network architectures with a constant global batch size of 4,096.
[…]
Compression ratios for AB trained models against the traditional data parallel baseline trained models on ImageNet-2012 for two network architectures with a constant global batch size of 4,096.
5. Discussion
Table 1. Average interconnect traffic, in GB/s, on each processes. $N$ represents either the number of nodes for the constant global batch size experiments or the global batch size in thousands for the constant local batch size experiments.
Constant Local
Batch Size
Constant Global
Batch Size
$N$ Baseline AB Baseline AB
ViT 2 $4.67\pm 0.00$ $2.10\pm 0.07$
4 $4.77\pm 0.06$ $2.12\pm 0.01$ $4.81\pm 0.05$ $1.86\pm 0.39$
8 $4.58\pm 0.00$ $1.76\pm 0.29$ $7.05\pm 1.03$ $3.45\pm 0.18$
16 $4.44\pm 0.10$ $1.74\pm 0.07$ $13.66\pm 0.16$ $5.33\pm 0.70$
32 $4.07\pm 0.20$ $1.58\pm 0.24$ $14.09\pm 3.47$ $7.88\pm 0.02$
ResNet-50 2 $2.06\pm 0.01$ $1.02\pm 0.00$
4 $2.07\pm 0.01$ $0.98\pm 0.05$ $2.06\pm 0.01$ $1.00\pm 0.01$
8 $2.00\pm 0.05$ $1.05\pm 0.05$ $2.67\pm 0.05$ $1.16\pm 0.14$
16 $1.96\pm 0.12$ $1.15\pm 0.00$ $3.28\pm 0.25$ $1.39\pm 0.21$
32 $1.62\pm 0.03$ $1.05\pm 0.01$ $3.50\pm 1.00$ $1.76\pm 0.47$
@@ 5. Discussion
Across both scaling scenarios, we find a consistent reduction of approximately 50% on average network traffic, demonstrating the effectiveness of our AB training method in reducing communication overheads. Comprehensive results are listed in Table 1, which further highlights the significant bandwidth requirements for training even moderately sized neural networks, a challenge that would be exacerbated for larger, state-of-the-art models.
@@
In 13 of the 18 experiments AB training maintains competitive performance compared to the traditional baseline using the same hyperparameters, Figures 3 and 6. Without the independently trained groups, the performance is noticeably reduced, see Figure 3(a), while the compression values are similar, Figure 5.
[…]
@@
Across all experiments, we observed favorable compression ratios (Figures 7 and 5). For runs which reached the same or better loss value, the Vision Transformer’s compression ratios range from 1.89:1 to 2.54:1. ResNet-50’s compression was less impressive, although still respectable (ranging from 1.19:1 to 1.72:1 for outperforming runs). This suggests a potential interaction between model architecture and the effectiveness of our method.
@@
Table 2 demonstrates the effectiveness of AB training compared to existing low-rank and pruning methods. We present both a standard benchmark (ResNet-50 on ImageNet-2012) and an idealized scenario (VGG16 on CIFAR-10). The latter highlights the potential of compression techniques due to VGG16’s over-parameterization and CIFAR-10’s relative simplicity (Chang et al., 2023). AB training outperformed expectations in this setting, achieving a 44.14:1 compression ratio (reducing the model to 2.27% of its original size) with negligible accuracy loss. This surpasses the compression achieved by ICP (3.67% of original size) and ABCPrune (11.26%).
@@
To compare the communication required for training the other model compression methods shown in Table 2 with traditional DP, we estimate their maximum communication savings achievable. As all of these methods iteratively reduce the network size, we assume that the all start with the full-rank representation and remove parameters during training. In out experiments, we found that once the network began to compress the network, it removed much of the network quickly, then removed small amounts during the remaining training steps. However, we do not know the rate at which parameters are removed. Taking this into account, we assume that models train close to full rank for 25% of training then close to their final model state for 75% of training. We describe the estimated communication reduction (ECR) as
(4) $\text{ECR}=100\%-F-Lc$
@@
where $F$ and $L$ are the percentages of training spent near full-rank and the most-compressed network state respectively and $c$ is the final compression ratio expressed as a percentage.
AB training demonstrates the greatest communication reduction in the ResNet-50 benchmark with the ECR. In practice, we measured an approximate reduction of 50% for all use-cases. This discrepancy is caused by other communications needed during training outside of the model synchronization. With the measurement, we outperform all models except OIALR and DLRT, which reduce the number of trainable parameters by training the $\Sigma$ matrix from a network’s weight’s SVD. In the VGG16 scenario, ICP (Chang et al., 2023) and AB training without using individual groups show slightly higher extimated communication reductions than AB training itself. This suggests that AB training offers superior communication efficiency until models reach extremely high compression levels.
While promising, our method also reveals challenges. As expected, we observe generalization improvements in all experiments, these are most noticable in Figure 4. This is likely due to the low-rank representation in combination with the independently trained groups; both of which have regularizing effects (see ‘No Groups’ measurements in Figure 6). However, at larger scales, both loss and accuracy degrade significantly. Combined with the decreasing compression ratios seen with increasing global batch size (Figure 5), this suggests that the averaging of independently trained subgroups becomes detrimental at extreme scales.
Similar trends in the fixed global batch size scenario (Figure 6) suggest an incompatibility between the very large global batch size and the much smaller batches used during the independent training phases. This likely leads to overly large updates in the independent groups, causing them to diverge from each other. Although this divergence increases the regularization effect of independent group training, is reduces the accuracy of the network. Our compression results (Figure 7) support this hypothesis. As the regularization effect strengthens with scale, the divergences between independent groups occur primarily in the smaller, yet important, singular values. When these divergences occur, averaging them together will cause the divergent singular values to shrink relative to the other values.
These challenges likely stem from suboptimal model merging or overly aggressive learning rate schedules. Since it is common practice to adjust the learning rate based on global batch size in traditional DP training, a similar strategy may be required for our method.
Our findings highlight the need for further exploration across diverse models, datasets, and hardware configurations. Understanding the interplay between network bandwidth limitations, and our method’s potential benefits in real-world environments is crucial. Finally, the observed accuracy degradation at scale motivates research into improved update mechanisms, potentially exploring non-average update rules, mixing matrices, or loss-based weighted averaging schemes.
@@
Table 2. Comparison of low-rank and pruning methods for ResNet-50 (on ImageNet-2012) and VGG16 (on CIFAR10). ’Difference to Baseline’ indicates validation top 1 performance relative to the original full-rank model in each study, with positive values denoting better performance. AB training used a global batch size of 4,096 for ImageNet and 1,024 for CIFAR10, achieving maximum top 1 accuracies of 75.67% and 91.87% respectively. The estimated communication reduction (ECR) is defined by Equation 4. AB training’s ECR assumes independent groups do not utilize the compute system’s interconnect. OIALR’s and DLRT’s ECR uses the compression of the trainable parameters as they report.
@@
ResNet-50 ImageNet-2012 AB +1.55 % $1.39:1$ 73.29 %*
AB (no groups) -0.02 % $1.27:1$ 15.94 %
OIALR (Coquelin et al., 2024) -1.72 % $1.21:1$ 63.64 %
DLRT (Schotthöfer et al., 2022) -0.56 % $1.85:1$ 64.35 %
PP-1 (Singh et al., 2019) -0.20 % $2.26:1$ 41.81 %
CP (He et al., 2017) -1.40 % $2.00:1$ 37.5 %
SFP (He et al., 2018) -0.20 % $2.39:1$ 43.62 %
ThiNet (Luo et al., 2017) -1.50 % 2.71 : 1 47.32 %
ABCPrune (Lin et al., 2020a) -2.15 % 1.84 : 1 34.24 %
VGG16 CIFAR-10 AB -0.23 % 44.14 : 1 65.60 %*
AB (no groups) -0.89 % 36.63 : 1 72.95 %
OIALR (Coquelin et al., 2024) +0.10 % 3.70 : 1 64.59 %
DLRT (Schotthöfer et al., 2022) -1.89 % 1.79 : 1 16.88 %
ABCPrune (Lin et al., 2020a) +0.06 % 8.83 : 1 66.51 %
ICP (Chang et al., 2023) -0.31 % 27.25 : 1 72.25 %
6. Conclusion

### v2

@@ IV-A Computational Environment
IV-B Datasets and Models
We used the ImageNet-2012 dataset, which contains 1.2 million images, for our scaling experiments. Basic image augmentation was applied, including normalization, random resizing, random cropping, and a random flip. We trained two models on this dataset: ResNet-50 [33] and Vision Transformer B/16 (ViT) [34], chosen for their widespread use and distinct architectures. We also trained VGG16 on the CIFAR-10 dataset. Basic image augmentation was applied in the same way as for the ImageNet experiments. All models were trained with the AdamW [38] optimizer. Other training hyperparameters can be found in Appendix 1.
IV-C Hyperparameter Considerations
As with any method, hyperparameter tuning is essential to performance. Key hyperparameters (HPs) of AB training include the warmup and full-rank rebound phase durations, the number of AB training iterations, the AB decomposition rank-reduction parameter, and the frequency of SVD and synchronization steps. Additionally, the latter will influence the trade-offs between regularization and communication efficiency.
A hyperparameter search on CIFAR-100 [7] with a minified Vision Transformer model (patch size of eight instead of 16, six heads instead of twelve, depth of six instead of twelve) was conducted using propulate [39] to determine the ideal settings for AB-Training’s hyperparameters. Using Weights and Biases [40], we measures the importance of these parameters via a Bayesian analysis of the completed runs. We find that the most important of these was the numABStesps. Controlling how many steps to train in low rank was crucial. However, the correlation of this variable to the final top-1 accuracy was only 0.225, indicating that there is a sweet spot between too many and too few steps. The number of warmup steps was also found to be quite important, with short warmups performing worse. The search and our analysis of the results suggested the following guidelines which are used for all experiments:
@@ IV-C Hyperparameter Considerations
IV-D Constant Local Batch Size Scaling
@@
Figure 2 shows the highest top 1 accuracy achieved by each model during training. The compression values for the trained models are shown in Figure 4(a).
@@
Figure 2: Highest top 1 accuracy measurements for each training run on ImageNet-2012 for two network architectures with a constant local batch size of 256. Global batch sizes range from 2,048 to 32,768 in powers of 2. Error are plotted, though not always visible.
IV-E Constant Global Batch Size Scaling
In this experiment, we maintain a fixed global batch size while scaling compute resources. This contrasts with the standard approach of keeping local batch sizes constant, emphasizing the interconnect traffic reductions achieved by AB training. This will reduce the number of computations per process per forward-backward pass. With a fixed global batch size of 4,096, a four-node run would have a local batch size of 256, whereas an eight-node run would have a local batch size of 128.
This experiment serves to show the performance of AB training as the time required to process a local batch shrinks, as is the case when the accelerators on a system are replaced but the network infrastructure remains the same. It also provides insight into the ability of independent training groups to learn meaningful representations when the batch size used for independent training is much smaller than the global batch size. Without the independent groups, the accuracy of the trained network should be similar throughout scaling. Results for these runs are shown in Figures 3, 4(b) and 5.
@@ IV-E Constant Global Batch Size Scaling
Figure 3: Highest top 1 accuracy measurements for each training run on ImageNet-2012 for two network architectures with a constant global batch size of 4,096. Error are plotted, though not always visible.
(a) Constant local batch size of 256.
(b) Constant global batch size of 4,096.
@@
Figure 4: Compression ratios for AB training with and without groups and a traditional DP baseline on ImageNet-2012 for two network architectures
(a) Scaled interconnect traffic. The ‘Peak Bandwidth’ is that of each port of each node’s Infiniband card.
(b) Training wall-clock time.
Figure 5: Scaled interconnect traffic and job wall-clock time for the ViT B/16 trained on ImageNet-2012. Scaling is based on the average time required for the calculation and communication of gradients on 4 nodes.
V Discussion
@@
Table II shows that our AB training method consistently achieves an average reduction of 70.31% in scaled network traffic across the scaling tests. The scaled network traffic is calculated as the average network traffic across the entire job scaled by the percentage of time required by the backward step during training (measured as 53.84% on four nodes). The traffic reduction highlights AB training’s effectiveness in mitigating the DP training communication bottleneck, a key concern in distributed training, especially for large models in HPC environments. These results underscore the significant bandwidth requirements for training even moderately sized neural networks, a challenge further exacerbated by larger models.
@@
AB training maintains competitive accuracy compared to traditional DP training in 13 of the 18 scaling experiments (Figures 2 and 3), using the same hyperparameters and in a similar time frame. Removing the independently trained groups noticeably reduces accuracy (Figure 2(a)), while compression values remain similar (Figure 4(a)). This demonstrates that AB training’s performance benefits stem from both the low-rank representation and the independent subgroup training.
Our large-scale experiments reveal the interplay between communication efficiency and model accuracy. In our ViT experiments with 32 nodes and a constant global batch size (Figure 5), we observe the aforementioned communication bottleneck, where the training time no longer decreases despite additional compute resources. However, AB training continues to see a reduction in time to train due to its reduced communication demands, even with the additional overhead stemming from the computation of the SVD for all model weights. ResNet-50’s performance plateaus earlier due to its computation becoming the bottleneck, though AB training and the traditional DP baseline have similar accuracy and require roughly the same time to train. AB training maintains or reduces the training time as compared to PyTorch’s DDP for all scaling measurements (see Table II). This is most obvious in the measurements shown in Figure 5, .
@@
Across all experiments, we observed favorable compression ratios for AB training(Figures 4(b) and 4(a) and Table II). For runs which equaled or exceeded the baseline network accuracy, the Vision Transformer’s compression ratios range from 1.89 : 1 to 2.54 : 1. ResNet-50’s compression was less impressive, although still respectable (ranging from 1.19 : 1 to 1.72 : 1 for outperforming runs). This suggests a potential interaction between model architecture and the effectiveness of our method.
@@
Table I demonstrates AB training’s effectiveness compared to existing low-rank and pruning methods. We showcase two standard benchmarks for these methods: ResNet-50 on ImageNet-2012 and VGG16 on CIFAR-10. The latter highlights the potential of compression techniques due to VGG16’s over-parameterization and CIFAR-10’s relative simplicity [27].
AB training outperformed expectations on CIFAR10, achieving a 44.14 : 1 compression ratio (reducing the model to 2.27% of its original size) with negligible accuracy loss. This surpasses the compression achieved by ICP [27] (3.67% of original size) and ABCPrune [25] (11.26%). For the ResNet-50 benchmark, AB training was the only one of the listed methods to outperform the baseline.
@@
To compare the communication required for training the other model compression methods shown in Table I with traditional DP, we estimate their maximum communication savings achievable. As all of these methods iteratively reduce the network size, we assume that they all start with the full-rank representation and remove parameters during training. In our experiments, we found that once the network began to compress the network, it removed much of the network quickly, then removed small amounts during the remaining training steps. However, we do not know the rate at which parameters are removed. Taking this into account, we assume that models train close to full rank for 25% of training then close to their final model state for 75% of training. We describe the estimated communication reduction (ECR) as
$\text{ECR}=100\%-F-Lc$ (4)
@@
where $F$ and $L$ are the percentages of training spent near full-rank and the most-compressed network state respectively and $c$ is the final compression ratio expressed as a percentage. For AB training, the percentage of training spent near the final compression state, $L$, is reduced by the percentage time spent in the independent group training phase.
@@
TABLE I: Comparison of low-rank and pruning methods for ResNet-50 (on ImageNet-2012) and VGG16 (on CIFAR10). ’Difference to Baseline’ indicates validation top 1 performance relative to the original full-rank model in each study, with positive values denoting better performance. AB training used a global batch size of 4,096 for ImageNet and 1,024 for CIFAR10, achieving maximum top 1 accuracies of 75.67% and 91.87% respectively. The estimated communication reduction (ECR) is defined by Equation 4. AB training’s ECR assumes independent groups do not utilize the compute system’s interconnect.
@@
ResNet-50 ImageNet-2012 AB +1.55 % $1.39:1$ 73.29 %
AB - No Groups -0.02 % $1.27:1$ 15.94 %
OIALR [5] -1.72 % $1.21:1$ 63.64 %
DLRT [11] -0.56 % $1.85:1$ 64.35 %
PP-1 [22] -0.20 % $2.26:1$ 41.81 %
CP [26] -1.40 % $2.00:1$ 37.5 %
SFP [23] -0.20 % $2.39:1$ 43.62 %
ThiNet [24] -1.50 % 2.71 : 1 47.32 %
ABCPrune [25] -2.15 % 1.84 : 1 34.24 %
VGG16 CIFAR-10 AB -0.23 % 44.14 : 1 65.60 %
AB - No Groups -0.89 % 36.63 : 1 72.95 %
OIALR [5] +0.10 % 3.70 : 1 64.59 %
DLRT [11] -1.89 % 1.79 : 1 16.88 %
ABCPrune [25] +0.06 % 8.83 : 1 66.51 %
ICP [27] -0.31 % 27.25 : 1 72.25 %
AB training demonstrates the greatest communication reduction in the ResNet-50 benchmark with the ECR. In practice, we measured an approximate network traffic reduction of 70% for all use-cases, landing quite close to the ECR. With the ECR, we outperform all models except OIALR and DLRT, which reduce the number of trainable parameters by training the $\Sigma$ matrix from a network’s weight’s SVD. In the VGG16 scenario, ICP [27] and AB training without using individual groups show slightly higher estimated communication reductions than AB training itself. This suggests that AB training offers superior communication efficiency until models reach extremely high compression levels.
Our results also reveal a complex interplay between low-rank representations, large batch sizes, and generalization. We observe significant generalization improvement in all experiments, likely due to the regularization effect of low-rank representations and independent group training (see ‘No Groups’ measurements in Figure 3). However, at larger scales, accuracy degrades. The decreasing compression ratios with increasing batch size in constant local batch size experiments (Figure 4(a)) suggest this degradation might be due to the loss or underutilization of important smaller singular values during model averaging.
Similarly, in constant global batch size experiments, increasing compression is coupled with decreasing accuracy (Figures 4(b) and 3). This is attributed to larger divergences in smaller singular values as local groups explore the loss landscape independently. As this effect strengthens, the divergences between independent groups occur primarily in the smaller, yet important, singular values. Averaging them together causes the divergent singular values to shrink relative to the other values. Over the course of training, these values can become so small that they are removed, hindering AB training at extreme scales.
[…]
Constant Local Batch Size
ViT B/16 ResNet-50
Batch Size AB AB - No Groups Traditional DDP AB AB - No Groups Traditional DDP
2k Scaled Traffic, $\mathrm{G}\mathrm{B}\mathrm{/}\mathrm{s}$ 2.81 $\pm$ 0.08 5.14 $\pm$ 0.19 8.87 $\pm$ 0.00 1.11 $\pm$ 0.00 2.87 $\pm$ 0.01 3.83 $\pm$ 0.01
Compression 2.52 : 1 2.83 : 1 1 : 1 1.29 : 1 1.27 : 1 1 : 1
Time to Train, $\min$ 435.80 $\pm$ 7.36 442.94 $\pm$ 8.52 436.38 $\pm$ 0.35 368.15 $\pm$ 0.80 364.91 $\pm$ 0.53 367.36 $\pm$ 0.36
Validation Top 1, % 70.24 $\pm$ 0.36 68.31 $\pm$ 0.07 70.73 $\pm$ 0.44 75.67 $\pm$ 0.10 74.42 $\pm$ 0.04 74.43 $\pm$ 0.06
4k Scaled Traffic, $\mathrm{G}\mathrm{B}\mathrm{/}\mathrm{s}$ 2.83 $\pm$ 0.06 5.72 $\pm$ 0.08 8.87 $\pm$ 0.10 1.04 $\pm$ 0.05 2.67 $\pm$ 0.04 3.85 $\pm$ 0.01
Compression 2.09 : 1 2.22 : 1 1 : 1 1.39 : 1 1.27 : 1 1 : 1
Time to Train, $\min$ 220.85 $\pm$ 0.88 224.35 $\pm$ 1.22 249.03 $\pm$ 31.01 200.69 $\pm$ 10.30 195.46 $\pm$ 1.30 195.19 $\pm$ 0.59
Validation Top 1, % 70.90 $\pm$ 0.44 67.81 $\pm$ 0.49 69.15 $\pm$ 0.00 74.61 $\pm$ 0.06 73.68 $\pm$ 0.12 73.67 $\pm$ 0.05
8k Scaled Traffic, $\mathrm{G}\mathrm{B}\mathrm{/}\mathrm{s}$ 2.65 $\pm$ 0.04 5.93 $\pm$ 0.13 8.51 $\pm$ 0.00 0.91 $\pm$ 0.03 2.71 $\pm$ 0.02 3.71 $\pm$ 0.09
Compression 2.02 : 1 1.98 : 1 1 : 1 1.31 : 1 1.16 : 1 1 : 1
Time to Train, $\min$ 115.65 $\pm$ 1.37 121.95 $\pm$ 1.95 121.78 $\pm$ 1.02 108.19 $\pm$ 4.94 104.70 $\pm$ 0.52 105.77 $\pm$ 1.69
Validation Top 1, % 68.92 $\pm$ 0.10 65.70 $\pm$ 0.14 67.20 $\pm$ 0.07 73.66 $\pm$ 0.03 73.04 $\pm$ 0.15 73.02 $\pm$ 0.10
16k Scaled Traffic, $\mathrm{G}\mathrm{B}\mathrm{/}\mathrm{s}$ 2.54 $\pm$ 0.03 4.91 $\pm$ 0.67 8.25 $\pm$ 0.18 0.87 $\pm$ 0.01 2.56 $\pm$ 0.04 3.63 $\pm$ 0.23
Compression 1.86 : 1 1.86 : 1 1 : 1 1.19 : 1 1.10 : 1 1 : 1
Time to Train, $\min$ 64.83 $\pm$ 1.96 79.15 $\pm$ 9.31 63.53 $\pm$ 1.10 58.61 $\pm$ 0.07 59.54 $\pm$ 0.68 59.27 $\pm$ 0.10
Validation Top 1, % 64.20 $\pm$ 0.01 61.17 $\pm$ 0.33 63.68 $\pm$ 0.09 72.66 $\pm$ 0.04 72.60 $\pm$ 0.10 72.32 $\pm$ 0.03
32k Scaled Traffic, $\mathrm{G}\mathrm{B}\mathrm{/}\mathrm{s}$ 2.27 $\pm$ 0.41 4.63 $\pm$ 0.51 7.52 $\pm$ 0.36 0.78 $\pm$ 0.01 2.17 $\pm$ 0.04 3.00 $\pm$ 0.06
Compression 1.80 : 1 1.84 : 1 1 : 1 1.20 : 1 1.12 : 1 1 : 1
Time to Train, $\min$ 39.53 $\pm$ 5.29 43.36 $\pm$ 3.90 36.56 $\pm$ 0.19 36.91 $\pm$ 0.07 37.48 $\pm$ 0.15 37.91 $\pm$ 0.73
Validation Top 1, % 55.54 $\pm$ 0.34 53.89 $\pm$ 0.15 58.00 $\pm$ 0.17 70.49 $\pm$ 0.05 71.38 $\pm$ 0.07 71.47 $\pm$ 0.08
Constant Global Batch Size
ViT B/16 ResNet-50
Nodes AB AB - No Groups Traditional DDP AB AB - No Groups Traditional DDP
4 Scaled Traffic, $\mathrm{G}\mathrm{B}\mathrm{/}\mathrm{s}$ 2.49 $\pm$ 0.49 5.61 $\pm$ 0.16 8.93 $\pm$ 0.09 1.07 $\pm$ 0.01 2.69 $\pm$ 0.03 3.83 $\pm$ 0.01
Compression 2.09 : 1 2.21 : 1 1 : 1 1.38 : 1 1.27 : 1 1 : 1
Time to Train, $\min$ 259.44 $\pm$ 49.68 227.99 $\pm$ 5.29 224.97 $\pm$ 1.79 194.64 $\pm$ 0.91 195.76 $\pm$ 3.49 195.03 $\pm$ 0.74
Validation Top 1, % 70.68 $\pm$ 0.56 68.66 $\pm$ 0.22 69.26 $\pm$ 0.18 74.75 $\pm$ 0.09 73.65 $\pm$ 0.13 73.74 $\pm$ 0.13
8 Scaled Traffic, $\mathrm{G}\mathrm{B}\mathrm{/}\mathrm{s}$ 4.89 $\pm$ 0.44 10.31 $\pm$ 0.32 13.09 $\pm$ 1.92 1.46 $\pm$ 0.18 3.98 $\pm$ 0.41 4.96 $\pm$ 0.09
Compression 2.27 : 1 2.19 : 1 1 : 1 1.54 : 1 1.27 : 1 1 : 1
Time to Train, $\min$ 123.35 $\pm$ 5.81 124.85 $\pm$ 1.92 153.03 $\pm$ 17.77 135.98 $\pm$ 12.43 131.76 $\pm$ 12.52 139.82 $\pm$ 2.00
Validation Top 1, % 69.66 $\pm$ 0.44 68.37 $\pm$ 0.15 69.38 $\pm$ 0.14 73.76 $\pm$ 0.19 73.60 $\pm$ 0.02 73.64 $\pm$ 0.09
16 Scaled Traffic, $\mathrm{G}\mathrm{B}\mathrm{/}\mathrm{s}$ 8.17 $\pm$ 0.17 17.14 $\pm$ 0.28 25.00 $\pm$ 0.31 1.91 $\pm$ 0.27 5.57 $\pm$ 0.09 6.09 $\pm$ 0.46
Compression 2.54 : 1 2.20 : 1 1 : 1 1.72 : 1 1.27 : 1 1 : 1
Time to Train, $\min$ 77.68 $\pm$ 9.18 74.26 $\pm$ 0.76 76.95 $\pm$ 0.87 111.46 $\pm$ 12.41 92.52 $\pm$ 1.78 113.00 $\pm$ 7.23
Validation Top 1, % 69.19 $\pm$ 0.68 68.67 $\pm$ 0.34 68.79 $\pm$ 0.40 72.41 $\pm$ 0.04 73.54 $\pm$ 0.04 73.72 $\pm$ 0.03
32 Scaled Traffic, $\mathrm{G}\mathrm{B}\mathrm{/}\mathrm{s}$ 9.84 $\pm$ 0.07 21.81 $\pm$ 2.66 25.00 $\pm$ 2.55 1.84 $\pm$ 0.42 5.70 $\pm$ 0.03 6.49 $\pm$ 1.85
Compression 2.91 : 1 2.20 : 1 1 : 1 1.91 : 1 1.27 : 1 1 : 1
Time to Train, $\min$ 51.10 $\pm$ 0.29 81.41 $\pm$ 19.36 77.99 $\pm$ 18.29 110.09 $\pm$ 28.61 92.37 $\pm$ 3.35 121.55 $\pm$ 31.04
Validation Top 1, % 66.13 $\pm$ 0.22 68.58 $\pm$ 0.57 68.99 $\pm$ 0.10 70.59 $\pm$ 0.08 73.73 $\pm$ 0.10 73.71 $\pm$ 0.04
Understanding the interplay between network bandwidth limitations and our method’s potential benefits in real-world environments is crucial. The observed accuracy degradation at scale suggests the need to research improved update mechanisms, potentially exploring non-average update rules, mixing matrices, or loss-based weighted averaging schemes.
VI Conclusion